# NumeroLogic: Number Encoding for Enhanced LLMs’ Numerical Reasoning

### v1

@@ Abstract
Figure 1: Reading numbers in a causal manner from left to right is sub-optimal for LLMs, as it is for humans. The model has to reach the final digits of the number before it can infer the place value of the first digit. To address this issue we propose “NumeroLogic”, a numerical format where the digit count is indicated before the actual number. Image adapted from imagery generated by DALL-E 3 [2].
@@ 1 Introduction
Large Language Models (LLMs) struggle with numerical and arithmetical tasks. Despite continuous improvements, even the most advanced models like GPT-4 [1] still exhibit poor performance when confronted with tasks such as multiplying 3-digit numbers [13]. Recent studies ([10, 13]) have proposed techniques to improve arithmetic in LLMs, such as the Chain of Thought (CoT; [16]) method, which pushes the model to anticipate the entire sequence of algorithmic steps rather than just the final output. While these strategies offer valuable insights into the capabilities of LLMs, they primarily concentrate on post-hoc solutions for specific arithmetic challenges and do not present a practical solution for pretraining LLMs. Our research, however, focuses on solutions applicable to self-supervised language modeling in general, utilizing arithmetic exercises primarily for evaluating their impact.
We hypothesize that one of the challenges LLMs face when dealing with numerical tasks is the textual representation of numbers. In today’s most popular decoder-based LLMs, each token attends only to previous tokens. When a model “reads” a token representing a digit (or multiple digits) it cannot tell its place value, i.e. ‘1’ can represent 1 million, 1 thousand, or a single unit. Only when reaching the end of the number might the model update its representation of the previous digit tokens to be related to their real place value.
To address this issue, we propose a straightforward reformatting technique called ”NumeroLogic,” which involves adding the number of digits as a prefix to numbers. This lets the model know in advance what is the place value of a digit before it is read. This simple change also offers another benefit, when the model is generating a number it needs to first reason about what is going to be the number of digits. This acts as a Chain of Thought (CoT) [16], encouraging the model to perform some reasoning before it begins to predict digits. Implementing the suggested reformatting does not necessitate any alterations to the model’s architecture; it can be accomplished through text pre- and post-processing based on regular expressions.
@@ 2 Related Work
Recently, there has been a significant interest in enhancing the numerical capabilities of LLMs. One approach to investigating these capabilities is by assessing their performance in arithmetic tasks. Several recent studies have proposed methods to enhance performance in these tasks. One strategy involves reversing the expected result order from the least significant digit to the most significant [10]. Another strategy is using an elaborated CoT where the model is taught to predict all steps of an algorithm predefined for each arithmetic task [10]. In [13], it is noted that the model learns to rely too heavily on positional encoding when trained for a specific arithmetic task. They suggest ways to overcome it, e.g. adding random whitespaces in the middle of numbers. These studies aim to enhance the performance of arithmetic tasks by offering tailored solutions to the associated challenges. In contrast, our focus is on identifying solutions that benefit general language modeling rather than just arithmetic tasks, with arithmetic tasks being used solely for measuring improvements.
Another aspect important for LLMs numerical capabilities is the tokenization process. The commonly used Byte Pair Encoding (BPE) based methods [4, 12] for tokenization are based on the corpus distribution and can split a number to tokens in unintuitive ways. Different foundation models took different approaches when dealing with number tokenization. PaLM [3], Llama [15], and Mistral [8] force each digit to have a single token. GPT-3.5 and GPT-4 define a token for each up to 3-digit number [1]. Somewhat related to our work, in [14], they highlighted an issue with the GPT approach. They show that dividing large numbers into 3-digit segments from left to right undermines arithmetic performance. They suggest overcoming it by artificially inserting commas between digits to control the splitting.
@@ 3 NumeroLogic
text = re.sub(pattern, , text)
text = re.sub(’<en>’, , text)
@@
For small transformers (NanoGPT [9]), we train all parameters from scratch with character-level tokenization. For small transformers, we also replace the special tokens with single characters, "<sn>", "<mn>", and "<en>" are replaced with "{", ":", and "}", respectively. For larger transformers, we start from pre-trained models. We add the new special tokens to the tokenizer’s vocabulary and expand the embedding layer and the final fully connected layer to fit the new vocabulary size. When continuing training on causal language modeling or fine-tuning on supervised arithmetic tasks, we use low-rank adaptation (LoRA) [7]. We apply LoRA for the attention block projection matrices (Q, K, V, O) and train the modified embedding layer and the final fully-connected layer in full rank.
@@ 4.1 Arithmetic tasks with small model
We trained NanoGPT [9] from scratch in a supervised manner jointly on 5 arithmetic tasks: addition, subtraction, multiplication, sine, and square root. Addition and subtraction are performed with up to 3-digit integer operands. Multiplications are performed with up to 2-digit integer operands. Sine and square root with 4 decimal-places floating point operands and results. The operand range for sine is within $\left[-\pi/2,\pi/2\right]$. The operand range for the square root is within $\left[0,10\right]$. The model is trained in a multi-task fashion on all 5 tasks, with 10K training samples for each task except for multiplication, for which 3K samples are used. We followed the protocol from Section D.2 in [10].
Figure 1 compares the results of training with plain numbers to training with the NumeroLogic encoding. For addition and subtraction, a model trained with plain numbers reached $88.37\%$ and $73.76\%$ accuracy, respectively, while with the NumeroLogic encoding, the tasks are almost solved ($99.96\%$ and $97.2\%$). For multiplication, we observe more than doubling of the accuracy, from $13.81\%$ to $28.94\%$. Furthermore, for the floating point operations, sine and square root, we see a significant improvement of ${}4\%$ for both tasks.
@@ 4.2 Arithmetic tasks with larger model
Next, we test how the method scales to a larger model. For this experiment, we fine-tune a pretrained Llama2-7B model [15]. In this experiment, we again tested the same five arithmetic tasks: addition, subtraction, multiplication, sine, and square root. For addition (5 digit), subtraction (5 digit), and multiplication (3 digit) we tested on two versions - integers and floating point numbers. For generating a random N-digit floating point operand we first sample an up to N-digit integer and then divide it by a denominator uniformly sampled from $\left\{10^{0},10^{1},...,10^{N}\right\}$. For each of the addition, subtraction, and multiplication tasks, we generated 300K random equations as a training set. The sine and square root operands and results are generated with 5 decimal place accuracy, we generated 30K random equations for the training sets of these tasks. Since we are working with a pretrained model we add new tokens ("<sn>", "<mn>", and "<en>") to the tokenizer’ vocabulary. We finetune one model per task with LoRA [7] (rank 8), we also train in full-rank the embedding layer and the final fully-connected layer since their parameters are extended to accommodate the larger vocabulary size.
The results are presented in Table 2. Addition and subtraction of integers are mostly solved by a model as large as Llama2-7B even for much larger numbers (e.g. 20-digit). For our 5-digit experiments, the plain text baselines reached 99.86% and 99.6% performance, for addition and subtraction, respectively. Despite the high performance of plain text, we still observe an improvement when using NumerLogic, with a perfect 100% for addition and rectification of more than 80% of the subtraction mistakes, reaching 99.93% accuracy for subtraction. For all other, non-saturated, tasks we observed significant gains of 1%-6%.
@@ 4.3 Self-Supervised Pretraining
Our approach differs from other methods in that it is not specialized for a specific task, such as arithmetic, but rather designed for general language modeling tasks involving text with numerical values. To test this capability we continue the pretraining of LLama2-7B with the causal text modeling objective (next token prediction). We train on text from the RefinedWeb dataset [11]. The goal is to teach the model to read and write numbers in the NumeroLogic format without forgetting its previously acquired knowledge. To facilitate this, we perform the continued pretraining with LoRA. We then test the model in a zero-shot manner on the massive multitask language understanding benchmark (MMLU) [6, 5].
In Figure 2, we present the MMLU 0-shot results obtained from training the model using plain numbers versus NumeroLogic encoding on an equal number of tokens. While training with plain numbers does not enhance the model’s accuracy compared to the pretrained model, employing NumeroLogic encoding results in a statistically significant improvement of 0.5%. The MMLU benchmark encompasses tasks from diverse domains, some emphasizing analytical skills and numerical comprehension while others do not. In Table 3, we delve into the impact of NumeroLogic on MMLU tasks categorized by field. As anticipated, tasks in STEM fields exhibit more substantial enhancements compared to those in social sciences and humanities. Table 4 provides a detailed analysis of NumeroLogic’s performance boost across tasks containing numbers versus those that do not. Consistently, tasks involving numbers show a more pronounced improvement.
@@
Table 6: Different encoding alternatives
4.4 Ablation studies
4.4.1 Encoding operands vs. results
We experimented to test the effect of operand encoding vs. the expected output (equation result) encoding. Operand encoding primarily influences the model’s comprehension of numerical values in the input, while result encoding is more associated with CoT, prompting the model first reason about the expected number of digits. We repeat the experiment from Section 4.1, but with the NumeroLogic encoding applied only to the operands or to the results and report the 3-digit addition results for the different variants. The results are presented in Table 5. We find that both operands and results encodings are beneficial, with a stronger impact attributed to encoding the results. Applying NumeroLogic to all numbers, both operands and results, yields the highest level of accuracy.
4.4.2 Different Encodings
We experimented with different formats for providing the number of digits. One alternative we tested is defining a set of new special tokens representing each possible number of digits, {<1digitnumber>, <2digitnumber>,...}. We observed that the performance of having multiple special tokens is even lower than plain numbers. It might be due to the unbalanced distribution of numbers. E.g. numbers with a single digit are much less frequent in a data of 3-digit additions, it is possible the model has not seen enough single-digit numbers to learn a good representation of the <1digitnumber> token. Another alternative we tested is removing the “end of number” token (<en>), keeping only the number prefix, e.g. "<sn>3<mn>100". This works better than plain but slightly worse than the full NumeroLogic encoding. The results are summarized in Table 6.
@@ 5 Conclusions
In this paper, we introduced NumeroLogic, an innovative approach to enhancing language models’ comprehension and generation of numerical data. Our hypothesis centered around the challenge of numerical representation in text for large language models (LLMs), which traditionally struggle with arithmetic and numerical understanding. By presenting numbers with a prefixed notation indicating their digit count, we proposed a solution that aids models in recognizing the place value of digits before the complete presentation of the full number. This method not only facilitates a better understanding of numbers but also prompts the model to reason about the magnitude of numbers it intends to generate, effectively integrating a form of Chain of Thought (CoT) reasoning.
Our experiments spanned from arithmetic tasks to general natural language understanding, using both small and large model architectures. The results consistently demonstrated that NumeroLogic significantly improves models’ numerical capabilities. In arithmetic tasks, we observed marked improvements in accuracy across different tasks including integer and floating-point operations. This improvement was not just confined to specialized arithmetic tasks but also extended to general language modeling, as evidenced by performance boosts in the MMLU benchmark for tasks requiring numerical understanding.
In conclusion, NumeroLogic offers a simple yet powerful tool for enhancing the numerical capabilities of language models without necessitating architectural modifications. Its success across various tasks and models highlights its potential as a general-purpose solution, particularly beneficial in domains where numerical understanding is crucial.

### v2

@@ 1 Introduction
Large Language Models (LLMs) struggle with numerical and arithmetical tasks. Despite continuous improvements, even the most advanced models like GPT-4 Achiam et al. (2023) still exhibit poor performance when confronted with tasks such as multiplying 3-digit numbers Shen et al. (2023). Recent studies (Lee et al. (2024); Shen et al. (2023)) have proposed techniques to improve arithmetic in LLMs, such as the Chain of Thought (CoT; Wei et al. (2022)) method, which pushes the model to anticipate the entire sequence of algorithmic steps rather than just the final output. While these strategies offer valuable insights into the capabilities of LLMs, they primarily concentrate on post-hoc solutions for specific arithmetic challenges and do not present a practical solution for pretraining LLMs. Our research, however, focuses on solutions applicable to self-supervised language modeling in general, utilizing arithmetic exercises primarily for evaluating their impact.
We hypothesize that one of the challenges LLMs face when dealing with numerical tasks is the textual representation of numbers. In today’s most popular decoder-based LLMs, each token attends only to previous tokens. When a model “reads" a token representing a digit (or multiple digits) it cannot tell its place value, i.e. ‘1’ can represent 1 million, 1 thousand, or a single unit. Only when reaching the end of the number might the model update its representation of the previous digit tokens to be related to their real place value.
Figure 1: Reading numbers in a causal manner from left to right is sub-optimal for LLMs, as it is for humans. The model has to reach the final digits of a number before it can infer the place value of the first digit. To address this, we propose “NumeroLogic", a numerical format where digit count is indicated before the actual number. Image by DALL-E 3 Betker et al. (2023).
To address this issue, we propose a straightforward reformatting technique called "NumeroLogic," which involves adding the number of digits as a prefix to numbers. This lets the model know in advance what is the place value of a digit before it is read. This simple change also offers another benefit, when the model is generating a number it needs to first reason about what is going to be the number of digits. This acts as a Chain of Thought (CoT) Wei et al. (2022), encouraging the model to perform some reasoning before it begins to predict digits. Implementing the suggested reformatting does not necessitate any alterations to the model’s architecture; it can be accomplished through text pre- and post-processing based on regex.
@@ 2 Related Work
Recently, there has been a significant interest in enhancing the numerical capabilities of LLMs. One approach to investigating these capabilities is by assessing their performance in arithmetic tasks. Several recent studies have proposed methods to enhance performance in these tasks. One strategy involves reversing the expected result order from the least to the most significant digit Lee et al. (2024). Another strategy is using an elaborated CoT where the model is taught to predict all steps of an algorithm predefined for each arithmetic task Lee et al. (2024). In Shen et al. (2023), it is noted that the model learns to rely too heavily on positional encoding when trained for a specific arithmetic task. They suggest ways to overcome it, e.g. adding random white spaces in the middle of numbers. These studies aim to enhance the performance of arithmetic tasks by offering tailored solutions to the associated challenges. In contrast, our focus is on identifying solutions that benefit general language modeling rather than just arithmetic tasks, with arithmetic tasks being used solely for measuring improvements.
Another aspect important for LLMs numerical capabilities is the tokenization process. The commonly used Byte Pair Encoding (BPE) based methods Gage (1994); Sennrich et al. (2015) for tokenization are based on the corpus distribution and can split a number to tokens in unintuitive ways. Different foundation models took different approaches when dealing with number tokenization. PaLM Chowdhery et al. (2023), Llama Touvron et al. (2023), and Mistral Jiang et al. (2023) force each digit to have a single token. GPT-3.5 and GPT-4 define a token for each up to 3-digit number Achiam et al. (2023). Somewhat related to our work, in Singh and Strouse (2024), they highlighted an issue with the GPT approach. They show that dividing large numbers into 3-digit segments from left to right undermines arithmetic performance. They suggest overcoming it by inserting commas between digits to control the splitting. Another related work, is Kim et al. (2021). They focus on the extrapolation ability of LLMs to unseen numbers and use a special number encoding that lets the LLM know the digit place-value.
@@ 3 NumeroLogic
For small transformers, we train all parameters from scratch with character-level tokenization. For small transformers, we also replace the special tokens with single characters, "<sn>", "<mn>", and "<en>" are replaced with "{", ":", and "}", respectively. For larger transformers, we start from pre-trained models. We add the new special tokens to the tokenizer’s vocabulary and expand the embedding layer and the final fully connected layer to fit the new vocabulary size. When continuing training on causal language modeling or fine-tuning on supervised arithmetic tasks, we use low-rank adaptation (LoRA) Hu et al. (2021). We apply LoRA for the attention block projection matrices (Q, K, V, O) and train the modified embedding layer and the final fully-connected layer in full rank.
@@
text = re.sub(pattern, ’’, text)
text = re.sub(’<en>’, ’’, text)
@@ 4.1 Arithmetic tasks with small model
We trained NanoGPT Karpathy (2022) from scratch in a supervised manner jointly on 5 arithmetic tasks: addition, subtraction, multiplication, sine, and square root. Addition and subtraction are performed with up to 3-digit integer operands. Multiplications are performed with up to 2-digit integer operands. Sine and square root with 4 decimal-places floating point operands and results. The operand range for sine is within $\left[-\pi/2,\pi/2\right]$. The operand range for the square root is within $\left[0,10\right]$. The model is trained in a multi-task fashion on all 5 tasks, with 10K training samples for each task except for multiplication, for which 3K samples are used. We followed the protocol from Section D.2 in Lee et al. (2024).
Tab. 1 compares the results of training with plain numbers to training with the NumeroLogic encoding. For addition and subtraction, a model trained with plain numbers reached $88.37\%$ and $73.76\%$ accuracy, respectively, while with the NumeroLogic encoding, the tasks are almost solved ($99.96\%$ and $97.2\%$). For multiplication, we observe more than doubling of the accuracy, from $13.81\%$ to $28.94\%$. Furthermore, for the floating point operations, sine and square root, we see a significant improvement of ${}4\%$ for both tasks.
@@ 4.2 Arithmetic tasks with larger model
Next, we test how the method scales to a larger model. For this experiment, we fine-tune a pretrained Llama2-7B model Touvron et al. (2023). In this experiment, we again tested the same five arithmetic tasks: addition, subtraction, multiplication, sine, and square root. For addition (5 digit), subtraction (5 digit), and multiplication (3 digit) we tested on two versions - integers and floating point numbers. For generating a random N-digit floating point operand we first sample an up to N-digit integer and then divide it by a denominator uniformly sampled from $\left\{10^{0},10^{1},...,10^{N}\right\}$. For each of the addition, subtraction, and multiplication tasks, we generated 300K random equations as a training set. The sine and square root operands and results are generated with 5 decimal place accuracy, we generated 30K random equations for the training sets of these tasks. Since we are working with a pretrained model we add new tokens ("<sn>", "<mn>", and "<en>") to the tokenizer’s vocabulary. We finetune one model per task with LoRA Hu et al. (2021) (rank 8), we also train in full-rank the embedding layer and the final linear layer since they are extended to fit the larger vocab. size.
The results are presented in Tab. 2. Addition and subtraction of integers are mostly solved by a model as large as Llama2-7B even for much larger numbers (e.g. 20-digit). For our 5-digit experiments, the plain text baselines reached 99.86% and 99.6% performance, for addition and subtraction, respectively. Despite the high performance of plain text, we still observe an improvement when using NumerLogic, with a perfect 100% for addition and rectification of more than 80% of the subtraction mistakes, reaching 99.93% accuracy for subtraction. For all other, non-saturated, tasks we observed significant gains of 1%-6%.
@@ 4.3 Self-Supervised Pretraining
Our approach differs from other methods in that it is not specialized for a specific task, such as arithmetic, but rather designed for general language modeling tasks involving text with numerical values. To test this capability we continue the pretraining of LLama2-7B with the causal text modeling objective (next token prediction). We train on text from the RefinedWeb dataset Penedo et al. (2023). The goal is to teach the model to read and write numbers in the NumeroLogic format without forgetting its previously acquired knowledge. To facilitate this, we perform the continued pretraining with LoRA. We then test the model in a 0-shot manner on MMLU Hendrycks et al. (2021b, a).
In Fig. 2, we present the MMLU 0-shot results obtained from training the model using plain numbers versus NumeroLogic encoding on an equal number of tokens. While training with plain numbers does not enhance the model’s accuracy compared to the pretrained model, employing NumeroLogic encoding results in a statistically significant improvement of 0.5%. The MMLU benchmark encompasses tasks from diverse domains, some emphasizing analytical skills and numerical comprehension while others do not. In Tab. 3, we delve into the impact of NumeroLogic on MMLU tasks categorized by field. As anticipated, tasks in STEM fields exhibit more substantial enhancements compared to those in social sciences and humanities. Tab. 4 provides a detailed analysis of NumeroLogic’s performance boost across tasks containing numbers versus those that do not. Consistently, tasks involving numbers show higher improvement.
@@
4.4 Ablation studies
4.4.1 Encoding operands vs. results
We experimented to test the effect of operand encoding vs. the expected output (equation result) encoding. Operand encoding primarily influences the model’s comprehension of numerical values in the input, while result encoding is more associated with CoT, prompting the model first reason about the expected number of digits. We repeat the experiment from Section 4.1, but with the NumeroLogic encoding applied only to the operands or to the results and report the 3-digit addition results for the different variants. The results are presented in Tab. 5. We find that both operands and results encodings are beneficial, with a stronger impact attributed to encoding the results. Applying NumeroLogic to all numbers, both operands and results, yields the highest level of accuracy.
4.4.2 Different Encodings
We experimented with different formats for providing the number of digits. One alternative we tested is defining a set of new special tokens representing each possible number of digits, {<1digitnumber>, <2digitnumber>,...}. We observed that the performance of having multiple special tokens is even lower than plain numbers. It might be due to the unbalanced distribution of numbers. E.g. numbers with a single digit are much less frequent in the data of 3-digit additions, it is possible the model has not seen enough single-digit numbers to learn a good representation of the <1digitnumber> token. Another alternative we tested is removing the “end of number" token (<en>), keeping only the number prefix, e.g. "<sn>3<mn>100". This works better than plain but slightly worse than the full NumeroLogic encoding. The results are summarized in Tab. 6.
@@ 4.4.2 Different Encodings
Table 6: Different encoding alternatives performance on 3-digit integer multiplications.
4.4.3 Is it the extra tokens?
It has been shown that the advantage of CoT is at least partially due to the extra tokens that allow the model to perform more computations or store information Pfau et al. (2024). To understand the effect of the extra tokens we run an experiment where all the extra tokens introduced by NumeroLogic are replaced with filler white-space tokens. Additionally, in Shen et al. (2023), it has been shown that the model learns to rely too heavily on positional encoding when trained on arithmetic tasks. It causes failures when the model is tested with numbers less frequent in the training data (e.g. 1 1-digit numbers when the model is trained on up to 3-digit numbers). To deal with this limitation, they suggest adding filler white-space tokens at random locations between the digits. We also report the results of their approach Shen et al. (2023) where we use the same number of tokens as NumeroLogic would have required, just that they are replaced with white-space tokens at random locations. These experiments were performed by finetuning Llama2-7B on 3-digit floating-point multiplication. The results are reported in Table 7. We observe that just adding the extra tokens does not help and the performance is similar to the plain format. Adding the same amount of extra tokens in random locations is somewhat helpful but not as effective as NumeroLogic. It eliminates the model’s reliance on positional encoding but does not provide place-value information like NumeroLogic.
Format Example Accuracy
NumeroLogic {1:1}*{1:1}={1:1} 31.03%
White-spaces ___1_*___1_=___1_ 24.37%
Random white-spaces ____1*__1__=1____ 27.76%
Plain 1*1=1 24.73%
Table 7: Extra tokens effect: Just adding filler white space tokens is not helpful and is comparable to the plain format. The random white-space method Shen et al. (2023) of adding filler tokens at random locations is helpful but less effective compared to NumeroLogic.
@@ 5 Conclusions
We introduced NumeroLogic, a novel method to improve language models’ handling of numerical data. Our approach prefixes numbers with their digit count, enhancing models’ understanding of place value and prompting better reasoning about numbers’ magnitude, akin to chain-of-thought reasoning. We tested NumeroLogic on both arithmetic and broader language understanding tasks. The results showed substantial enhancements in numerical tasks, including integer and floating-point calculations, and in broader modeling contexts like the MMLU benchmark. In summary, NumeroLogic is a straightforward yet effective enhancement for language models’ numerical abilities, applicable across various tasks without requiring changes to the models’ architecture.
6 Limitations
The NumeroLogic encoding, while enhancing numerical reasoning, might increase the number of tokens per number. Moreover, it introduces additional steps in pre- and post-processing. This raises the computational costs and also potentially increases the model’s latency during inference. These factors might impact the efficiency of NumeroLogic, especially in numerical-processing-intensive applications.
Our experiments predominantly involved fine-tuning pre-trained language models (LLMs) rather than training them from scratch with NumeroLogic. While this limits our ability to conclusively predict the impacts from the pre-training phase, incorporating NumeroLogic early in the pre-training would likely have a stronger positive rather than negative effect on the performance. Additionally, our testing did not extend to models larger than 7B parameters. However, it has been demonstrated that both small and large models exhibit similar learning behaviors Warstadt et al. (2023); therefore, it is plausible to predict that scaling up the model size will not diminish the effectiveness of NumeroLogic.
Lastly, our evaluation was confined to controlled academic benchmarks, which might not fully represent the complexities of real-world numerical data. Extending testing to diverse, real-world datasets is essential to fully understand NumeroLogic’s practical effectiveness and ensure it can handle the unpredictable nature of real-world numerical data. Similarly, despite caring mainly about numerical aspects, we checked English-focused datasets and data. The cross effects with different languages, scripts and even numerical writing system is left as an open question.